OpenStudy (anonymous):

let f(x)=x+8 and g(x)=x^2-6x-7 find f(x)+g(x)

OpenStudy (campbell_st):

well add f and g and collect like terms \[f(x) + g(x) = x + 8 + x^2 -6x - 7\] which is \[f(x) + g(x) = x^2 + x - 6x + 8 - 7\] collect the like terms and simplify for the answer.